OpenStudy (anonymous):

I really need this explained to me. The sum of two consecutive even integers is 234. What is the larger integer?

OpenStudy (anonymous):

An even integer is one that can be divided by 2.

OpenStudy (anonymous):

2x is always even.

jimthompson5910 (jim_thompson5910):

consecutive even integers are integers like: 2, 4, 6, 8, 10, 12, 14 they are whole numbers that are even that are right next to each other let x be the first integer the next consecutive even integer is x+2 add the two up to get x+(x+2) x+x+2 2x+2 then set this equal to 234 to get 2x+2 = 234 now solve for x

OpenStudy (anonymous):

@jim_thompson5910 x=115?

jimthompson5910 (jim_thompson5910):

2x+2 = 234 2x+2-2 = 234-2 2x = 232 x = ???

jimthompson5910 (jim_thompson5910):

x is even by the way

OpenStudy (anonymous):

x=116?

jimthompson5910 (jim_thompson5910):

so if x = 116, then x+2 = ???

OpenStudy (anonymous):

x+2=118

jimthompson5910 (jim_thompson5910):

so 118 is your final answer

OpenStudy (anonymous):

@jim_thompson5910 thanks for all your help!

jimthompson5910 (jim_thompson5910):

you're welcome